# Spiral renormalization group flow and universal entanglement spectrum of the non-Hermitian 5-state Potts model

### v2

@@ Abstract
Tensor networks—We define the ground state wavefunction of the deformed Potts Hamiltonian as being the eigenstate corresponding to the eigenvalue with the smallest real part, and assume that the Schmidt coefficients decay fast enough such that it can be represented as a matrix product state (MPS) [33]. In particular, we impose the $\mathbb{Z}_{5}$ subgroup symmetry of the Hamiltonian for computational efficiency and to identify the charges of the CCFT primary operators easily. We use an MPS ansatz with open BC (and a Hamiltonian with periodic BC, hence with one long-ranged interaction term) such that can employ the standard density matrix renormalization group (DMRG) [35]. The excited states are determined by an algorithm that targets the quasiparticle excitations [32, 12]. We find ground states with a bond dimension $\chi=400$, while for excited states we use $\chi=600$, both with an algorithmic tolerance of $10^{-6}$. We simulate up to system sizes $L=64$, assuring that we are only in the finite-size scaling regime and that second-order terms remain negligible [22]. Due to the small non-Hermitian deformation, the singular values of the MPS retain their ordering, explaining why the MPS representations obtained through DMRG and the Quasiparticle Ansatz (QPA) remain robust. As further analyzed in [26], the algorithms only experience slight systematic errors in the imaginary results, a known problem [27]. We use the Julia packages TensorKit.jl and MPSKit.jl to perform the numerical simulations [9, 31].
@@
We now extract the conformal data with the refined critical parameter. To do so, we fit Eq. (S.5) to the fit parameters $\Delta_{n}$, $f$, $c$, $a$ and $\tilde{a}_{n}$ for L=814L=8-14. The conformal scaling dimension estimates are visible in Table 1. We compare these values with the theoretical predictions found in [10]. Additionally, we include the scaling dimensions extracted through exact diagonalization (ED) as done in [30] for $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and L=813L=8-13. We observe that the improved accuracy of our new fixed point yields more accurate estimates for the conformal dimensions. The ED result at $\lambda_{c}^{\text{prev}}$ only has better accuracy compared to the MPS algorithms at $\lambda_{c}$ for the imaginary parts of the conformal data. This is most likely due to the non-Hermitian MPS ansatz having more untrustworthy low entanglement contributions [26]. Moreover, the tensor network framework deteriorates with increasing scaling dimension due to deviations from the area law. Lastly, the imaginary part of $\Delta_{\varepsilon}$ in Table 1 was fitted for $1/L^{2}$, which according to Fig. 2 does not appear to be correct. The scaling which is closest to the theoretical value is $1/L^{1.05}$ with a value of $\mathrm{Im}(\Delta_{\varepsilon})=-0.2245(1)i$. It remains unclear whether this observation has a deeper meaning or if it is due to the same simulation artifacts discussed above.
@@
Fig. 3 shows the theoretical scaling dimension of these operators and their degeneracies. We find that the logarithm of the eigenvalues of the half-chain density matrix $\rho^{\mathcal{A}}_{RL}$ indeed yields the entanglement spectrum. The entanglement spectrum largely agrees with the free-free boundary CFT spectrum. For example, we identify the degeneracies of $\phi_{5,1}$ with its scaling dimension $\Delta_{\phi_{5,1}}=4-0.9190i$, as there are three indistinguishable levels in the zero sector and two in the one charged sector. Only the imaginary part of the entanglement spectrum seems to converge more slowly. This could be caused not only by MPS limitations [26], but also because $\rho_{RL}$ is not positive semi-definite, making the relation between the entanglement Hamiltonian and the density matrix less apparent. As observed in [27], only a slight systematic error in the imaginary parts of the achieved energies occurs when working with an MPS [26]. However, since the model has only a small non-Hermitian part, we still find that Eq. (7) yields information about the boundary CCFT. The entanglement spectrum on OBC will, although less explicitly, also converge to the free-free boundary CFT [26].
@@
The entanglement spectrum we determined through the entanglement Hamiltonian on open boundary conditions (OBC) is visible in Figure S5. Although it largely agrees with the PBC spectrum presented in the main text, inconsistencies also occur: the real part of the spectrum sometimes diverges from the predicted values, and the imaginary part exhibits what appears to be level crossings for $L\geq 32$. Moreover, we argue that the latter undergoes a potential branch cut in the logarithm, as it drastically jumps from positive to negative values without passing through any level crossings. We attribute these consistencies to the unreliability of ground state simulations in capturing the entanglement structure for OBC, as was also observed in the total entanglement scaling [6]. The OBC causes an abrupt end to the lattice, where entanglement cannot flow through, resulting in significant boundary effects. Lastly, the boundary spectrum was, for both PBC and OBC, nearly identical with the previous fixed point estimation $\lambda_{c}^{\text{prev}}=0.079+0.060i$.
@@
A large source of errors still arises from the fixed point accuracy. As discussed in Section G G, even though the precision of the fixed point was increased by an additional digit with respect to previous research [8], the repulsive nature of the point and its spiral RG flow punishes even small deviations from the true fixed point. Consequently, even for small deviations from the true fixed point, the energy levels will deviate from the expected scaling for large system sizes. We illustrate this in Figure S6, where the deviations from the theoretical scaling dimensions at $\lambda_{c}=0.0788+0.0603i$ are plotted for system sizes up to $L=64$. This repulsiveness is why the system sizes utilized in the predictions of the scaling dimensions are relatively small compared to what is computationally possible. This error is inherent to the nature of complex conformal field theories and can be resolved by repeating the procedure in Section G G for larger system sizes. However, as will be discussed next, simulation errors prevent this from being done reliably.
@@
Working within the MPS manifold leads to different sources of simulation errors. First, the MPS representation has a finite bond dimension. This introduces deviations from the true wavefunction because only finite entanglement is present. As a result, there is a perturbation from the true critical point, preventing additional precision of the fixed point [Huang2024]. Moreover, for critical systems, where the ground state’s correlation length diverges, and for excited states, which may violate the area law of entanglement, this approximation can be particularly problematic. The limited entanglement at the MPS boundary also introduces additional boundary effects. However, we confirmed that DMRG and QPA reach their bond saturation within the algorithmic tolerance for $\chi\geq 400$ and $600$, respectively, for system sizes up to $L=24$, regardless of the additional difficulties associated with simulating non-Hermitian models.

### v3

@@ Abstract
Tensor networks—We define the ground state wavefunction of the deformed Potts Hamiltonian as being the eigenstate corresponding to the eigenvalue with the smallest real part, and assume that the Schmidt coefficients decay fast enough such that it can be represented as a matrix product state (MPS) [33]. In particular, we impose the $\mathbb{Z}_{5}$ subgroup symmetry of the Hamiltonian for computational efficiency and to identify the charges of the CCFT primary operators easily. We use an MPS ansatz with open BC (and a Hamiltonian with periodic BC, hence with one long-ranged interaction term) such that we can employ the standard density matrix renormalization group (DMRG) [35]. The excited states are determined by an algorithm that targets the quasiparticle excitations [32, 12]. We find that the ground state and excitation energies deviate slightly from the values obtained by exact diagonalization (ED) as shown in Fig. 2. However, the errors decrease with increasing bond dimension with an algorithmic tolerance $10^{-6}$. This can be understood from the fact that the Hamiltonian (1) contains only a small portion of non-Hermiticity, thereby rendering the MPS truncation almost optimal. In the following, we use $\chi=400$ and $\chi=600$ for the ground state and excitation calculations. We also confirmed that the errors are in the same order as the algorithmic tolerance in the larger systems.
Figure 2: The difference between the ground state energy $E_{0}$ and a reference energy $E^{\text{ref}}_{0}$ at the critical point $\lambda_{c}$ with increasing bond dimension. For $L=12$ the reference is the ED ground state and for $L=24$ it is $E_{0}(\chi=450)$.
Consequently, we can achieve accurate results up to system sizes $L=24$, while ensuring that what remains are the finite-size effects and that second-order terms stay negligible [22]. In particular, we use larger system sizes beyond the reach of ED methods up to $L=64$ for computing the entanglement spectrum. Employing larger systems is crucial, for instance, in Fig. 5, where the maximal reach of ED methods $L\sim 12$ barely captures the right physics due to the logarithmic finite-size effects. The relatively good performance of the MPS algorithms on non-Hermitian models in capturing the entanglement can also be explained by the small non-Hermiticity. It retains the same ordering as the true entanglement structure. Consequently, as further analyzed in [26], the algorithms only experience slight systematic errors in the imaginary results, a known problem [27]. We use the Julia packages TensorKit.jl and MPSKit.jl to perform the numerical simulations [9, 31].
@@
We now extract the conformal data with the refined critical parameter. To do so, we fit Eq. (S.5) to the fit parameters $\Delta_{n}$, $f$, $c$, $a$ and $\tilde{a}_{n}$ for L=814L=8-14. The conformal scaling dimension estimates are visible in Table 1. We compare these values with the theoretical predictions found in [10]. Additionally, we include the scaling dimensions extracted through ED as done in [30] for $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and L=813L=8-13. We observe that the improved accuracy of our new fixed point yields more accurate estimates for the conformal dimensions. The ED result at $\lambda_{c}^{\text{prev}}$ only has better accuracy compared to the MPS algorithms at $\lambda_{c}$ for the imaginary parts of the conformal data. This is most likely due to the non-Hermitian MPS ansatz having more untrustworthy low entanglement contributions [26]. Moreover, the tensor network framework deteriorates with increasing scaling dimension due to deviations from the area law. Lastly, the imaginary part of $\Delta_{\varepsilon}$ in Table 1 was fitted for $1/L^{2}$, which according to Fig. 3 does not appear to be correct. The scaling which is closest to the theoretical value is $1/L^{1.05}$ with a value of $\mathrm{Im}(\Delta_{\varepsilon})=-0.2245(1)i$. It remains unclear whether this observation has a deeper meaning or if it is due to the same simulation artifacts discussed above.
@@
Figure 4: The fifteen largest eigenvalues $e_{i}$ of the $\mathbb{Z}_{5}$ charged and non-charged sectors of the reduced density matrix $\rho^{\mathcal{A}}_{RL}$ for a half-chain bipartition, normalized such that the first eigenvalue of each sector is one, converging for increasing bond dimension $\chi$ to the reference values $e_{i}(\chi=450)$.
@@
Fig. 5 shows the theoretical scaling dimension of these operators and their degeneracies. We find that the logarithm of the eigenvalues of the half-chain density matrix $\rho^{\mathcal{A}}_{RL}$ indeed yields the entanglement spectrum. The entanglement spectrum largely agrees with the free-free boundary CFT spectrum. For example, we identify the degeneracies of $\phi_{5,1}$ with its scaling dimension $\Delta_{\phi_{5,1}}=4-0.9190i$, as there are three indistinguishable levels in the zero sector and two in the one charged sector. Only the imaginary part of the entanglement spectrum seems to converge more slowly. This could be caused not only by MPS limitations [26], but also because $\rho_{RL}$ is not positive semi-definite, making the relation between the entanglement Hamiltonian and the density matrix less apparent. However, since both the scaling dimensions and the degeneracies are exactly as predicted in [29], we conclude that Eq. (7) yields information about the boundary CCFT. The entanglement spectrum on OBC will, although less explicitly, also converge to the free-free boundary CFT [26].
@@
The entanglement spectrum we determined through the entanglement Hamiltonian on open boundary conditions (OBC) is visible in Figure S5. Although it largely agrees with the PBC spectrum presented in the main text, inconsistencies also occur: the real part of the spectrum sometimes diverges from the predicted values, and the imaginary part exhibits what appears to be level crossings for $L\geq 32$. Moreover, we argue that the latter undergoes a potential branch cut in the logarithm, as it drastically jumps from positive to negative values without passing through any level crossings. We attribute these inconsistencies to the unreliability of ground state simulations in capturing the entanglement structure for OBC, as was also observed in the total entanglement scaling [6]. The OBC causes an abrupt end to the lattice, where entanglement cannot flow through, resulting in significant boundary effects. Lastly, the boundary spectrum was, for both PBC and OBC, nearly identical with the previous fixed point estimation $\lambda_{c}^{\text{prev}}=0.079+0.060i$.
@@
A large source of errors still arises from the fixed point accuracy. As discussed in Section G G, even though the precision of the fixed point was increased by an additional digit with respect to previous research [8], the repulsive nature of the point and its spiral RG flow punishes even small deviations from the true fixed point. Consequently, even for small deviations from the true fixed point, the energy levels will deviate from the expected scaling for large system sizes. We illustrate this in Figure S6, where the deviations from the theoretical scaling dimensions at $\lambda_{c}=0.0788+0.0603i$ are plotted for system sizes up to $L=64$. This repulsiveness is why the system sizes utilized in the predictions of the scaling dimensions are relatively small compared to what is computationally possible. This error is inherent to the nature of complex conformal field theories and can be resolved by repeating the procedure in Section G G for larger system sizes. However, as discussed next, simulation errors prevent this from being done reliably.
@@
Working within the MPS manifold leads to different sources of simulation errors. First, the MPS representation has a finite bond dimension. This introduces deviations from the true wavefunction because only finite entanglement is present. As a result, there is a perturbation from the true critical point, preventing additional precision of the fixed point [Huang2024]. Moreover, for critical systems, where the ground state’s correlation length diverges, and for excited states, which may violate the area law of entanglement, this approximation can be particularly problematic. The limited entanglement at the MPS boundary also introduces additional boundary effects. However, we confirmed that DMRG and QPA reach their bond saturation within (near) algorithmic tolerance for $\chi\geq 400$ and $600$, respectively, for system sizes up to $L=24$, regardless of the additional difficulties associated with simulating non-Hermitian models.